I need your help on proofs
SecX-TanX=(1-Sinx)/Cosx

Answer by stanbon(75887)

I need your help on proofs
SecX-TanX=(1-Sinx)/Cosx
-----
Convert everything to sin and cos
------
(1/cos) - (sin/cos) = (1-sin)/cos
-------
(1-sin)/cos = (1-sin)/cos
----
QED
---------
Cheers,
Stan H.
------